# Ultrastable 2D glasses and packings explained by local centrosymmetry

### v2

@@ Abstract
Using the most recent numerical data by Bolton-Lum et al. [Phys. Rev. Lett. 136, 058201 (2026)], we demonstrate that ideal ultrastable glasses in the athermal limit (or ultrastable ideal 2D disk packings) possess a remarkably high degree of local centrosymmetry. In particular, we find that the inversion-symmetry order parameter for local force transmission introduced in Milkus and Zaccone, [Phys. Rev. 93, 094204 (2016)], is as high as $F_{IS}=0.93546$, to be compared with $F_{IS}=1$ for perfect centrosymmetric crystals free of defects, and with $F_{IS}\sim 0.3-0.5$ for standard random packings. This observation provides a clear, natural explanation for the ultra-high shear modulus of ideal packings and ideal glasses, because the high centrosymmetry prevents non-affine relaxations which decrease the shear modulus. The same mechanism explains the absence of boson peak-like soft vibrational modes. These results also confirm what was found previous work, i.e. that the bond-orientational order parameter is a very poor correlator for the vibrational and mechanical properties of disordered packings.
@@
The central result of this work is the demonstration that ideal two-dimensional disk packings and ultrastable glasses possess an exceptionally high degree of local inversion symmetry, quantitatively captured by the inversion-symmetry order parameter $F_{\mathrm{IS}}$. The measured value $F_{\mathrm{IS}}=0.93546$ is remarkably close to the crystalline limit $F_{\mathrm{IS}}=1$ and stands in sharp contrast to the much smaller values ($F_{\mathrm{IS}}\sim 0.3-0.5$) characteristic of ordinary jammed packings and structural glasses. This finding provides a direct and physically transparent structural explanation for the long-standing puzzle of why ideal glasses exhibit crystal-like mechanical and vibrational properties despite the absence of long-range order. The large value of $F_{\mathrm{IS}}$ measured here naturally accounts for both the anomalously high rigidity and the crystal-like vibrational spectrum reported for ideal disk packings, without invoking hidden crystalline order or exotic mechanisms.
@@
More broadly, these results suggest that local centrosymmetry constitutes a key organizing principle for amorphous solids at the edge between glassy and crystalline behavior. Ideal glasses and ideal packings emerge as systems that remain topologically disordered yet are statistically optimized to maximize local inversion symmetry, thereby suppressing nonaffine softening. This insight offers a unifying geometric framework to understand ultrastability, hyperuniform mechanical response, and the disappearance of low-frequency vibrational excitations. Beyond two-dimensional disk packings, the same concepts are expected to apply to three-dimensional glasses, jammed materials, and metallic glasses, opening the door to symmetry-based design principles for mechanically robust amorphous materials.

### v3

@@ Abstract
Using the most recent numerical data by Bolton-Lum et al. [Phys. Rev. Lett. 136, 058201 (2026)], we demonstrate that ideal ultrastable glasses in the athermal limit (or ultrastable ideal 2D disk packings) possess a remarkably high degree of local centrosymmetry. In particular, we find that the inversion-symmetry order parameter for local force transmission introduced in Milkus and Zaccone, [Phys. Rev. 93, 094204 (2016)], is as high as $F_{IS}=0.84522$, to be compared with $F_{IS}=1$ for perfect centrosymmetric crystals free of defects, and with $F_{IS}\sim 0.3-0.5$ for standard random packings. This observation provides a clear, natural explanation for the ultra-high shear modulus of ideal packings and ideal glasses, because the high centrosymmetry prevents non-affine relaxations which decrease the shear modulus. The same mechanism explains the absence of boson peak-like soft vibrational modes. These results also confirm what was found previous work, i.e. that the bond-orientational order parameter is a very poor correlator for the vibrational and mechanical properties of disordered packings.
@@
In the above expressions, all summations over nearest neighbors are performed as directed sums $\sum_{i}\sum_{j\in\mathrm{nn}(i)}$, such that each bond contributes independently to the force imbalance of both particles it connects. In other words, each bond must contribute separately to each endpoint of the Gabriel graph, which is consistent with $F_{IS}$ being a network-averaged measure of local inversion symmetry directly interpretable as “how centrosymmetric the force environment of a typical node (atom $i$) is”.
Although the order parameter $F_{\mathrm{IS}}$ is constructed from the affine force field $\boldsymbol{\Xi}$, it is determined entirely by the static contact geometry and does not require evaluation of the dynamical (Hessian) matrix or elastic response. Therefore, the observed correlation between $F_{\mathrm{IS}}$ and the shear modulus is predictive rather than tautological, as $G$ depends additionally the Hessian and its exact calculation is much more demanding.
@@
The central result of this work is the demonstration that ideal two-dimensional disk packings and ultrastable glasses possess an exceptionally high degree of local inversion symmetry, quantitatively captured by the inversion-symmetry order parameter $F_{\mathrm{IS}}$. The measured value $F_{\mathrm{IS}}=0.84522$ is remarkably close to the crystalline limit $F_{\mathrm{IS}}=1$ and stands in sharp contrast to the much smaller values ($F_{\mathrm{IS}}\sim 0.3-0.5$) characteristic of ordinary jammed packings and structural glasses. This finding provides a direct and physically transparent structural explanation for the long-standing puzzle of why ideal glasses exhibit crystal-like mechanical and vibrational properties despite the absence of long-range order. The large value of $F_{\mathrm{IS}}$ measured here naturally accounts for both the anomalously high rigidity and the crystal-like vibrational spectrum reported for ideal disk packings, without invoking hidden crystalline order or exotic mechanisms.
@@
More broadly, these results suggest that local centrosymmetry constitutes a key organizing principle for amorphous solids at the edge between glassy and crystalline behavior. Ideal glasses and ideal packings emerge as systems that remain topologically disordered yet are statistically optimized to maximize local inversion symmetry, thereby suppressing nonaffine softening. This insight offers a unifying geometric framework to understand ultrastability, hyperuniform mechanical response, and the disappearance of low-frequency vibrational excitations. Beyond two-dimensional disk packings, the same concepts are expected to apply to three-dimensional glasses, jammed materials (including frictional suspensions and shear-thickening systems [DAmico2025]), and metallic glasses [Wei-Hua2025], opening the door to symmetry-based design principles for mechanically robust amorphous materials.